# Optimal Control of an Epidemic with Intervention Design

### v2

@@ Abstract
This paper investigates the optimal control of an epidemic governed by a SEIR model with operational delays in vaccination and non pharmaceutical interventions. We address the mathematical challenge of imposing hard healthcare capacity constraints (e.g., ICU limits) over an infinite time horizon. To rigorously bridge the gap between theoretical constraints and numerical tractability, we employ a variational framework based on Moreau–Yosida regularization and establish the connection between finite- and infinite-horizon solutions via $\Gamma$-convergence. The necessary conditions for optimality are derived using the Pontryagin Maximum Principle, allowing for the characterization of singular regimes where the optimal strategy maintains the infection level precisely at the capacity boundary. Numerical simulations illustrate these theoretical findings, quantifying the shadow prices of infection and costs associated with intervention delays.
@@ 1 Introduction
Second, we resolve the ambiguity surrounding healthcare capacity constraints. We formulate the problem with a "hard" state constraint $i(t)\leq I_{\max}$ to represent absolute ICU capacity limits. To facilitate numerical analysis without sacrificing theoretical rigor, we utilize Moreau–Yosida regularization to establish a formal link between this hard constraint and a penalized objective function Dal Maso (1993b); Rockafellar and Wets (1998). We prove that the solution to the penalized problem converges to the hard-constrained optimum as the penalty parameter increases, ensuring that our numerical results are theoretically consistent with strict infrastructure limits.
@@
Finally, we employ $\Gamma$-convergence to rigorously justify the transition from finite-horizon numerical simulations to infinite-horizon theoretical stability. Extending the recent results of Freddi and Goreac Freddi and Goreac (2024) and Della Rossa and Freddi Della Rossa and Freddi (2025) to the SEIR context with delays, we show that the optimal strategies obtained over large finite horizons approximate the true infinite-horizon optimum. This analysis allows us to characterize singular regimes—specifically, a "boundary-maintenance arc" where the system is optimally held at capacity—providing a meaningful characterization of the optimal strategy that descriptive models often miss.
@@
Notation: Throughout this paper, $I=[0,\infty)$ denotes the infinite time horizon. We utilize the space $L^{\infty}(I)$ for essentially bounded measurable functions and $W^{1,\infty}(I;\mathbb{R}^{3})$ for Lipschitz continuous state trajectories. The notation $u_{n}\rightharpoonup^{\ast}u$ denotes weak-* convergence in the control space. For a thorough mathematical treatment of these functional spaces, we refer the reader to Fonseca and Leoni (2007) and Leoni (2017).
@@ 2 Analysis of Controlled System
Here, $\beta$ is the baseline transmission rate, and the effective transmission rate $\beta-h(t)$ decreases as the intervention $h(t)$ increases.
@@
Let $I=[0,+\infty)$. Under Assumption 2.1, there exists a unique global solution $(s,e,i,r)\in W^{1,\infty}_{\mathrm{loc}}(I;\mathbb{R}^{4})$ to (1). Moreover, the trajectory remains in the compact simplex $\{(s,e,i,r)\in[0,1]^{4}:\ s+e+i+r=1\}$, so the vector field is uniformly bounded; in particular there exists $C>0$ such that $\|(s,e,i,r)^{\prime}\|_{L^{\infty}(I)}\leq C$. For all $t\geq 0$,
@@
$\tau:=\inf\{t\geq 0:\ e(t)<0\ \text{or}\ i(t)<0\},\qquad\inf\emptyset:=+\infty.$
@@
    From (1), $s^{\prime}(t)=-[(\beta-h(t))i(t)+u(t)]s(t)$. We have established $s(t)>0$ for all $t\geq 0$ and $i(t)>0$ for every finite $t\geq 0$ (Theorem 2.2).
    Also, Assumption 2.1 guarantees $\beta-h(t)\geq\beta-h_{\max}>0$. Since $u(t)\geq 0$, the term in brackets is strictly positive. Thus, $s^{\prime}(t)<0$ for all $t$.
@@ 3.1 Contribution and Framework
This section resolves this ambiguity through a variational framework that explicitly links the two formulations. We begin by establishing the existence of optimal controls for a penalized infinite-horizon problem with discounting and bounded controls. Crucially, we introduce a theoretical bridge based on Yosida regularization, demonstrating that the quadratic over-capacity penalty term is the Moreau–Yosida envelope of the hard constraint indicator. This result proves that the optimal solutions of the penalized problem converge to the hard-constrained optimum as the penalty parameter approaches infinity Rockafellar and Wets (1998). Furthermore, we provide an explicit topological proof of $\Gamma$-convergence to justify the use of finite-horizon approximations for the infinite-horizon problem Dal Maso (1993a). Finally, we characterize the optimality conditions via the Pontryagin Maximum Principle (PMP) applied to the penalized problem. We identify singular regimes where the switching function vanishes, showing that the planner optimally maintains the epidemic near a dynamic boundary rather than applying purely bang–bang controls. Throughout the analysis, we utilize the controlled SEIR dynamics (System 1) under Assumption 2.1, working with the reduced state vector $x(t):=(s(t),e(t),i(t))$ since $r(t)$ is redundant by population conservation.
@@ 3.2 Problem Formulations
Let $I=[0,\infty)$. The reduced state is $x(t)=(s(t),e(t),i(t))\in[0,1]^{3}$, and the control is $v(t)=(u(t),h(t))\in[0,u_{\max}]\times[0,h_{\max}]$. The control $u(t)$ represents the vaccination rate (moving mass from $s$ to $r$), while $h(t)$ represents suppression effort (reducing transmission from $\beta$ to $\beta-h(t)$). Distinct from both is the healthcare capacity constraint $i(t)\leq I_{\max}$.
@@ 3.3.1 Topology and Convergence Framework
$\int_{0}^{t}u_{n}s_{n}-\int_{0}^{t}us=\int_{0}^{t}u_{n}(s_{n}-s)+\int_{0}^{t}(u_{n}-u)s,$
the first term converges to 0 because $\|u_{n}\|_{\infty}$ is uniformly bounded and $s_{n}\to s$ in $L^{1}$, while the second term converges to 0 by weak-$\ast$ convergence of $u_{n}$ tested against the fixed $L^{1}$ function $s\,\mathbf{1}_{[0,t]}$. Similarly,
$\int_{0}^{t}h_{n}(s_{n}i_{n})-\int_{0}^{t}h(si)=\int_{0}^{t}h_{n}(s_{n}i_{n}-si)+\int_{0}^{t}(h_{n}-h)(si),$
@@
    Compactness: The set $\mathcal{U}_{\mathrm{ad}}$ is a subset of the ball $\{v\in L^{\infty}(I;\mathbb{R}^{2}):\|v\|_{\infty}\leq M\}$ where $M=\max(u_{\max},h_{\max})$. This ball is weak-* compact by Banach–Alaoglu. Since $L^{1}(I)$ is separable, the weak-* topology on this ball is metrizable, implying weak-* sequential compactness. The box constraints and the linear delay constraint are weak-* closed (for the delay: take test functions supported on $0,t_{\text{delay}})$, then $\int u_{n}\varphi=0$ implies $\int u\varphi=0$, so $u=0$ a.e. on the delay interval). Thus, $\mathcal{U}_{\mathrm{ad}}$ is weak-* sequentially compact. There exists a subsequence (not relabeled) such that $(u_{n},h_{n})\rightharpoonup^{\ast}(u^{\kappa},h^{\kappa})\in\mathcal{U}_{\mathrm{ad}}$.
@@
    Term-by-Term Convergence: We show $J_{\kappa}(u^{\kappa},h^{\kappa})\leq\liminf_{n}J_{\kappa}(u_{n},h_{n})$. The integrand involves bilinear terms ($us$, ihih) and the penalty. For any finite $M>0$, consider the term $\int_{0}^{M}u_{n}(t)s_{n}(t)e^{-\delta t}dt$. We decompose it as:
    $\int_{0}^{M}u_{n}(s_{n}-s^{\kappa})e^{-\delta t}dt+\int_{0}^{M}(u_{n}-u^{\kappa})s^{\kappa}e^{-\delta t}dt+\int_{0}^{M}u^{\kappa}s^{\kappa}e^{-\delta t}dt.$
    The first integral vanishes because $u_{n}$ is bounded and $s_{n}\to s^{\kappa}$ uniformly. The second vanishes because $(u_{n}-u^{\kappa})\rightharpoonup^{\ast}0$ and $s^{\kappa}e^{-\delta t}\in L^{1}([0,M])$. The same decomposition applies to $i_{n}h_{n}$ and $i_{n}$. Additionally, $\psi(i_{n})\to\psi(i^{\kappa})$ uniformly on $[0,M]$, ensuring convergence of the penalty integral. Thus, the integrals converge on $[0,M]$. Finally, since $i(t)\in[0,1]$, we have $0\leq\psi(i(t))\leq(1-I_{\max})^{2}$, hence the integrand is bounded by $C_{\kappa}e^{-\delta t}\in L^{1}([0,\infty))$, and the tail contribution vanishes as $M\to\infty$. (See also Dacorogna (2007) for general existence results via direct methods).
@@ 3.3.2 Moreau–Yosida regularization: linking soft and hard constraints
We now rigorously link the two problems. The hard constraint set $C$ can be represented by the extended-value indicator function $\iota_{C}(x)$. Pointwise in the scalar variable $i$, the Moreau–Yosida envelope of the scalar indicator $\iota_{(-\infty,I_{\max}]}(i)$ with parameter $\varepsilon>0$ is defined as:
@@
2. Recovery sequence: Let $(u,h)\in\mathcal{F}$. Consider the constant sequence $(u_{n},h_{n})=(u,h)$. Since the trajectory is feasible, $\psi(i(t;u,h))\equiv 0$. Thus, $J_{\kappa_{n}}(u,h)=J_{\mathrm{strict}}(u,h)$ for all $n$. By the Fundamental Theorem of $\Gamma$-convergence Dal Maso (1993a), minimizers converge. ∎
@@ 3.4 Variational convergence: finite to infinite horizon
2. Recovery sequence: Take the constant sequence $u_{n}=u,h_{n}=h$. The difference $|J_{\kappa}-J_{\kappa,T_{n}}|$ is the tail integral $\int_{T_{n}}^{\infty}\ell_{\kappa}dt$. Since the state space is compact and controls are bounded, $\ell_{\kappa}(t,x,v)\leq Ce^{-\delta t}\in L^{1}$. By the Dominated Convergence Theorem, the tail integral vanishes as $n\to\infty$. ∎
@@ 3.5 Optimality conditions and economic interpretation
We apply Pontryagin’s Maximum Principle (PMP) to the finite-horizon penalized problem $(\mathcal{P}_{\kappa,T})$ on $[0,T]$, with free terminal state (implying the transversality condition $\lambda(T)=0$). Define the Hamiltonian:
@@
Since $\psi(i)=({i-I_{\max}})_{+}^{2}$ is continuously differentiable, the adjoint equations hold in the classical sense (see Mangasarian (1966) for sufficient conditions and Avram et al. (2021) for PMP in similar contexts).

### v3

@@ Abstract
This paper investigates the optimal control of an epidemic governed by a SEIR model with an operational delay in vaccination. We address the mathematical challenge of imposing hard healthcare capacity constraints (e.g., ICU limits) over an infinite time horizon. To rigorously bridge the gap between theoretical constraints and numerical tractability, we employ a variational framework based on Moreau–Yosida regularization and establish the connection between finite- and infinite-horizon solutions via $\Gamma$-convergence. The necessary conditions for optimality are derived using the Pontryagin Maximum Principle, allowing for the characterization of boundary-maintenance arcs where the optimal strategy maintains the infection level precisely at the capacity boundary. Numerical simulations illustrate these theoretical findings, quantifying the shadow prices of infection and costs associated with intervention delays.
@@ 1 Introduction
Second, we resolve the ambiguity surrounding healthcare capacity constraints. We formulate the problem with a "hard" state constraint $i(t)\leq I_{\max}$ to represent absolute ICU capacity limits. To facilitate numerical analysis without sacrificing theoretical rigor, we utilize Moreau–Yosida regularization to establish a formal link between this hard constraint and a penalized objective function Dal Maso (1993); Rockafellar and Wets (1998). We prove that the solution to the penalized problem converges to the hard-constrained optimum as the penalty parameter increases, ensuring that our numerical results are theoretically consistent with strict infrastructure limits.
@@
Finally, we employ $\Gamma$-convergence to rigorously justify the transition from finite-horizon numerical simulations to infinite-horizon theoretical stability. Extending the recent results of Freddi and Goreac Freddi and Goreac (2024) and Della Rossa and Freddi Della Rossa and Freddi (2025) to the SEIR context with delays, we show that the optimal strategies obtained over large finite horizons approximate the true infinite-horizon optimum. This analysis allows us to characterize boundary-maintenance arcs, where the system is optimally held at capacity—providing a meaningful characterization of the optimal strategy that descriptive models often miss.
@@
Notation: Throughout this paper, $I=[0,\infty)$ denotes the infinite time horizon. We utilize the space $L^{\infty}(I)$ for essentially bounded measurable functions and $W^{1,\infty}(I;\mathbb{R}^{k})$ for Lipschitz continuous state trajectories. The subscript $\mathrm{loc}$” denotes the local variant: $f\in W^{1,\infty}_{\mathrm{loc}}(I;\mathbb{R}^{k})$ if $f\in W^{1,\infty}([0,T];\mathbb{R}^{k})$ for every $T>0$. The notation $u_{n}\rightharpoonup^{\ast}u$ denotes weak-* convergence in the control space. For a thorough mathematical treatment of these functional spaces, we refer the reader to Fonseca and Leoni (2007) and Leoni (2017).
@@ 2 Analysis of Controlled System
Here, $\beta$ is the baseline transmission rate, and the effective transmission rate $(\beta-h(t))$ decreases as the intervention $h(t)$ increases.
@@
Let $I=[0,+\infty)$. Under Assumption 2.1, there exists a unique global solution $(s,e,i,r)\in W^{1,\infty}_{\mathrm{loc}}(I;\mathbb{R}^{4})$ to (1). Moreover, the trajectory remains in the compact simplex $\{(s,e,i,r)\in[0,1]^{4}:\ s+e+i+r=1\}$, so the vector field is uniformly bounded; in particular there exists $K>0$ such that $\|(s,e,i,r)^{\prime}\|_{L^{\infty}(I)}=\|(\dot{s},\dot{e},\dot{i},\dot{r})\|_{L^{\infty}(I)}\leq K$. For all $t\geq 0$,
@@
$\tau:=\inf\{t\geq 0:\ e(t)<0\ \text{or}\ i(t)<0\},\qquad\inf\emptyset:=+\infty\quad$
where the last convention means that if the set is empty—i.e., neither $e(t)$ nor $i(t)$ ever becomes negative—then $\tau=+\infty$ and the trajectory remains non-negative for all time.
@@
Remark 2.3 (Role of the vaccination delay).
The delay parameter $t_{\mathrm{delay}}$ constrains the vaccination control to $u(t)=0$ for $t<t_{\mathrm{delay}}$, reflecting the logistical lead time required for vaccine development or deployment. During this initial period, the planner must rely exclusively on the suppression instrument $h(t)$. The theoretical results of Sections 2 and 3 hold for any $t_{\mathrm{delay}}\geq 0$; the delay simply restricts the admissible set $\mathcal{U}_{\mathrm{ad}}$ without altering the structure of the existence or convergence proofs. Economically, a larger $t_{\mathrm{delay}}$ increases the shadow price of infection during the early phase, as the planner has fewer instruments available.
@@
    From (1), $s^{\prime}(t)=-[(\beta-h(t))i(t)+u(t)]s(t)$. We have established $s(t)>0$ for all $t\geq 0$ and $i(t)>0$ for every finite $t\geq 0$ (Theorem 2.2). Also, Assumption 2.1 guarantees $\beta-h(t)\geq\beta-h_{\max}>0$.
@@ 3.1 Contribution and Framework
The choice of an infinite planning horizon $I=[0,\infty)$ is motivated by both mathematical and practical considerations. From a practical standpoint, the terminal date of an epidemic is not known a priori; a finite-horizon formulation introduces an artificial terminal boundary that distorts the optimal policy near the end of the planning period through transversality effects. The infinite-horizon formulation, combined with discounting at rate $\delta>0$, eliminates this artifact and yields time-consistent policies whose structure does not depend on an arbitrarily chosen endpoint. This modeling choice is standard in the economics literature on dynamic optimization and resource management Kamien and Schwartz (1991); Acemoglu (2009); Sethi and Thompson (2000), and has been adopted in recent infinite-horizon epidemic control studies Freddi and Goreac (2024); Della Rossa and Freddi (2025); Bliman et al. (2021). In the epidemiological context, the infinite horizon is best understood not as a claim that the epidemic literally persists forever—indeed, Corollary 2.6 guarantees $i(t)\to 0$—but rather as a device ensuring that the optimal policy accounts for all future consequences without truncation bias.
This section resolves this ambiguity through a variational framework that explicitly links the two formulations. We begin by establishing the existence of optimal controls for a penalized infinite-horizon problem with discounting and bounded controls. Crucially, we introduce a theoretical bridge based on Yosida regularization, demonstrating that the quadratic over-capacity penalty term is the Moreau–Yosida envelope of the hard constraint indicator. This result proves that the optimal solutions of the penalized problem converge to the hard-constrained optimum as the penalty parameter approaches infinity Rockafellar and Wets (1998). Furthermore, we provide an explicit topological proof of $\Gamma$-convergence to justify the use of finite-horizon approximations for the infinite-horizon problem Dal Maso (1993). Finally, we characterize the optimality conditions via the Pontryagin Maximum Principle (PMP) applied to the penalized problem. We identify boundary-maintenance arcs where the switching function vanishes, showing that the planner optimally maintains the epidemic near a dynamic boundary rather than applying purely bang–bang controls. Throughout the analysis, we utilize the controlled SEIR dynamics (System 1) under Assumption 2.1, working with the reduced state vector $x(t):=(s(t),e(t),i(t))$ since $r(t)$ is redundant by population conservation.
@@ 3.2 Problem Formulations
Let $I=[0,\infty)$. The reduced state is $x(t)=(s(t),e(t),i(t))\in[0,1]^{3}$, and the control is $v(t)=(u(t),h(t))\in[0,u_{\max}]\times[0,h_{\max}]$. The control $u(t)$ represents the vaccination rate (moving mass from $s$ to $r$), while $h(t)$ represents suppression effort (reducing transmission from $\beta$ to $(\beta-h(t))$). Distinct from both is the healthcare capacity constraint $i(t)\leq I_{\max}$.
@@ 3.3.1 Topology and Convergence Framework
$\int_{0}^{t}u_{n}s_{n}\,d\xi-\int_{0}^{t}us\,d\xi=\int_{0}^{t}u_{n}(s_{n}-s)\,d\xi+\int_{0}^{t}(u_{n}-u)s\,d\xi,$
the first term converges to 0 because $\|u_{n}\|_{\infty}$ is uniformly bounded and $s_{n}\to s$ in $L^{1}$, while the second term converges to 0 by weak-$\ast$ convergence of $u_{n}$ tested against the fixed $L^{1}$ function $s\,\mathbf{1}_{[0,t]}$ (where $\mathbf{1}_{[0,t]}(\xi)=1$ if $\xi\in[0,t]$). Similarly,
$\int_{0}^{t}h_{n}(s_{n}i_{n})\,d\xi-\int_{0}^{t}h(si)\,d\xi=\int_{0}^{t}h_{n}(s_{n}i_{n}-si)\,d\xi+\int_{0}^{t}(h_{n}-h)(si)\,d\xi,$
@@
    Compactness: The set $\mathcal{U}_{\mathrm{ad}}$ is a subset of the ball $\{v\in L^{\infty}(I;\mathbb{R}^{2}):\|v\|_{\infty}\leq M\}$ where $M=\max\{u_{\max},h_{\max}\}$. This ball is weak-* compact by Banach–Alaoglu. Since $L^{1}(I)$ is separable, the weak-* topology on this ball is metrizable, implying weak-* sequential compactness. The box constraints and the linear delay constraint are weak-* closed (for the delay: for any test function $\varphi\in L^{1}(I)$ supported on $0,t_{\text{delay}})$, we have $\int u_{n}\varphi\,dt=0$ for all $n$, which implies $\int u\varphi\,dt=0$ by weak-$\ast$ convergence, so $u=0$ a.e. on the delay interval). Thus, $\mathcal{U}_{\mathrm{ad}}$ is weak-* sequentially compact. There exists a subsequence (not relabeled) such that $(u_{n},h_{n})\rightharpoonup^{\ast}(u^{\kappa},h^{\kappa})\in\mathcal{U}_{\mathrm{ad}}$.
@@
    Term-by-Term Convergence: We show $J_{\kappa}(u^{\kappa},h^{\kappa})\leq\liminf_{n}J_{\kappa}(u_{n},h_{n})$. The integrand involves bilinear terms ($us$, ihih) and the penalty. For any finite $\bar{T}>0$, consider the term $\int_{0}^{\bar{T}}u_{n}(t)s_{n}(t)e^{-\delta t}\,dt$. We decompose it as:
    $\int_{0}^{\bar{T}}u_{n}(s_{n}-s^{\kappa})e^{-\delta t}\,dt+\int_{0}^{\bar{T}}(u_{n}-u^{\kappa})s^{\kappa}e^{-\delta t}\,dt+\int_{0}^{\bar{T}}u^{\kappa}s^{\kappa}e^{-\delta t}\,dt.$
    The first integral vanishes because $u_{n}$ is bounded and $s_{n}\to s^{\kappa}$ uniformly. The second vanishes because $(u_{n}-u^{\kappa})\rightharpoonup^{\ast}0$ and $s^{\kappa}e^{-\delta t}\in L^{1}([0,\bar{T}])$. The same decomposition applies to $i_{n}h_{n}$ and $i_{n}$. Additionally, $\psi(i_{n})\to\psi(i^{\kappa})$ uniformly on $0,\bar{T}]$, ensuring convergence of the penalty integral. Thus, the integrals converge on $0,\bar{T}]$. Finally, since $i(t)\in[0,1]$, we have $0\leq\psi(i(t))\leq(1-I_{\max})^{2}$, hence the integrand is bounded by $C_{\kappa}e^{-\delta t}\in L^{1}([0,\infty))$, and the tail contribution vanishes as $\bar{T}\to\infty$. (See also Dacorogna (2007) for general existence results via direct methods).
@@ 3.3.2 Moreau–Yosida regularization: linking soft and hard constraints
We now rigorously link the two problems. The hard constraint set $C=\{x\in[0,1]^{3}:\ i\leq I_{\max}\}$, as defined in Section 3.2, can be represented by the extended-value indicator function $\iota_{C}(x)$. Pointwise in the scalar variable $i$, the Moreau–Yosida envelope of the scalar indicator $\iota_{(-\infty,I_{\max}]}(i)$ with parameter $\varepsilon>0$ is defined as:
@@
2. Recovery sequence:Let $(u,h)\in\mathcal{F}$ (the set of feasible controls defined in Section 3.2). Consider the constant sequence $(u_{n},h_{n})=(u,h)$. Since the trajectory is feasible, $\psi(i(t;u,h))\equiv 0$. Thus, $J_{\kappa_{n}}(u,h)=J_{\mathrm{strict}}(u,h)$ for all $n$. By the Fundamental Theorem of $\Gamma$-convergence Dal Maso (1993), minimizers converge. ∎
@@ 3.4 Variational convergence: finite to infinite horizon
2. Recovery sequence: Take the constant sequence $u_{n}=u,h_{n}=h$. The difference $|J_{\kappa}-J_{\kappa,T_{n}}|$ is the tail integral $\int_{T_{n}}^{\infty}\ell_{\kappa}\bigl(t,x(t;u,h),v(t)\bigr)\,dt$. Since the state space is compact and controls are bounded, $\ell_{\kappa}(t,x,v)\leq Ce^{-\delta t}\in L^{1}$. By the Dominated Convergence Theorem, the tail integral vanishes as $n\to\infty$. ∎
@@ 3.5 Optimality conditions and economic interpretation
We apply the Pontryagin Maximum Principle (PMP) to the finite-horizon penalized problem $(\mathcal{P}_{\kappa,T})$ on $[0,T]$, with free terminal state (implying the transversality condition $\lambda(T)=0$). Define the Hamiltonian:
@@
Since $\psi(i)=(\max\{0,\,i-I_{\max}\})^{2}$ is continuously differentiable, the adjoint equations hold in the classical sense (see Mangasarian (1966) for sufficient conditions, Avram et al. (2021) for PMP in similar contexts, and Aseev and Kryazhimskiy (2007); Carlson et al. (1991); Aseev and Kryazhimskiy (2004) for comprehensive treatments of the Pontryagin Maximum Principle on infinite-horizon problems).